A stationary body explodes into two fragments of masses m1 and m2. If momentum of one fragment is p, the energy of explosion is

Grade:12

2 Answers

Vikas TU
7 years ago
Momentum conserving,
0 = p + p’
p’ = -p
where p’ is the momentum of mass m2 say after exploding.
And Hence,
Energy,
E1 = p^2/2m1
and
E2 = p^2/2m2
Net energy
Enet = E1 + E2
       = p^2/2m1 + p^2/2m2
ANUSH NAYAK
6 years ago
 
Momentum conserving,
0 = p + p’
p’ = -p
where p’ is the momentum of mass m2 say after exploding.
And Hence,
Energy,
E1 = p^2/2m1
and
E2 = p^2/2m2
Net energy
Enet = E1 + E2
       = p^2/2m1 + p^2/2m2